### v1

# A hot spots theorem for the mixed eigenvalue problem with small Dirichet region

Abstract.
We prove that on convex domains, first mixed Laplace eigenfunctions have no interior critical points if the Dirichlet region is connected and sufficiently small. We use this result to construct a new family of polygonal domains for which Rauch’s hot spots conjecture holds and to prove a new general theorem regarding the hot spots conjecture.
1. Introduction
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a bounded, convex Lipschitz domain. Let $D\subseteq\overline{\Omega}$ (where $\overline{\Omega}$ is the closure of $\Omega$ in ${\mathbb{R}}^{2}$) be a compact set. Consider the eigenvalue problem
(1) $\begin{cases}-\Delta u=\lambda u\;\;&\text{in}\;\;\Omega\setminus D\\ u\equiv 0\;\;&\text{in}\;\;D\\ \partial_{\nu}u\equiv 0\;\;&\text{in}\;\;\partial\Omega\setminus D.\end{cases}$
This eigenvalue problem admits a discrete, increasing sequence of non-negative eigenvalues $\{\lambda_{n}^{D}\}$ and a corresponding orthonormal basis of eigenfunctions $\{u_{n}^{D}\}$ for $L^{2}(\Omega)$ that (weakly) solve (1). When we wish to emphasize the dependence on the domain, we may write $\lambda_{n}^{D}(\Omega):=\lambda_{n}^{D}$. Note that because we did not place strong restrictions on $D$, the (weak) eigenfunctions may not vanish everywhere on $D$ (see [FNO21]). We will suppose throughout the paper that $D$ is chosen such that eigenfunctions of (1) vanish everywhere on $D$. For example, this holds if $D$ is the closure of a relatively open subset of $\partial\Omega$ or if $D$ is the closure of a Lipschitz subdomain of $\Omega$.
For $\epsilon>0$, let $B(0,\epsilon)$ be the disk of radius $\epsilon$ centered at the origin in ${\mathbb{R}}^{2}$.
Theorem 1.1.
There exists $\epsilon>0$ such that if $D\subseteq B(0,\epsilon)$ and $D$ is connected, then $u_{1}^{D}$ has no critical points in $\Omega\setminus D$ [^1] Recall that a domain is, by definition, open, so the theorem does not prohibit critical points in $\partial\Omega$.
Remark 1.2.
If $D$ is compactly contained in $\Omega$, then the hypothesis that $D$ be connected is necessary. We demonstrate this in Example 4.1 below. If $D$ is contained in $\partial\Omega$, then it is unclear to us whether connectivity is a necessary condition or if its requirement is merely an artifact of the proof.
Remark 1.3.
Our proof of Theorem 1.1 is heavily inspired by Miyamoto’s proof of Theorem A in [M09].
The well-known hot spots conjecture of J. Rauch (see [R74]) states that extrema of solutions to the heat equation (with generic initial conditions) on a free membrane tend toward the boundary of the membrane as time tends toward infinity. It is not hard to see (see [BB99]) that this statement is equivalent to second Neumann eigenfunctions of the Laplace operator (i.e. second eigenfunctions of (1) with $D=\emptyset$) having extrema only on the boundary. A natural generalization of this problem is to ask for which sets $D$ does the heat equation with perfect refrigeration on $D$ and perfect insulation on $\partial\Omega\setminus D$ have solutions with extrema tending toward the boundary over time. Theorem 1.1 shows that if $\Omega$ is convex and $D$ is connected and sufficiently small, then this is the case.
In the past year or so, there has been an increased focus by researchers on variants of the hot spots conjecture for the mixed problem. In 2024 alone, the following pre-prints have appeared: [H24a], [LY24], [AR24], and [H24b]. To our knowledge these are the first papers on the topic since the early 2000s. See [AR24] for an overview of earlier results.
Theorem 1.1 is the combination of Propositions 1.5 and 2.1 below.
Definition 1.4.
Let $\Omega\subseteq{\mathbb{R}}^{2}$ be a simply connected, bounded Lipschitz domain. Let $D\subseteq\overline{\Omega}$ be as above. For each $y\in\partial\Omega$, let $\nu(y)$ be the outward normal vector to the boundary at $y$. We say that the pair $(\Omega,D)$ is Neumann convex if for any $x\in\Omega$ and $y\in\partial\Omega\setminus D$, the dot product $(y-x)\cdot\nu(y)$ is non-negative.
Note that $\Omega$ need not be convex for $(\Omega,D)$ to be Neumann convex. However, convexity of $\Omega$ implies Neumann convexity of $(\Omega,D)$ for any choice of $D$. Moreover, if $\Omega$ is a nodal domain of a second Neumann eigenfunction of a convex domain $\tilde{\Omega}$ and $D$ equals the nodal set of the eigenfunction, then $(\Omega,D)$ is Neumann convex.
Proposition 1.5.
Suppose that $(\Omega,D)$ is Neumann convex and that $D$ is connected. Let $d$ be the diameter of $\Omega$, and let $j_{0}$ be the first zero of the Bessel function $J_{0}(x)$. If
$\lambda_{1}^{D}\leq\Big{(}\frac{j_{0}}{d}\Big{)}^{2},$
then $u_{1}^{D}$ has no critical points in $\Omega\setminus D$.
In Section 4, we show some explicit examples where this inequality holds, and we use these results to prove two new results on the hot spots conjecture. Let $P_{n}$ denote a regular polygon with $n$ edges for $n\geq 3$. Let $e$ be the interior of an edge of $P_{n}$, and let $Q_{n}$ be the reflection of $P_{n}$ over $e$. Let $K_{n}=P_{n}\cup e\cup Q_{n}$. Then we have
Theorem 1.6.
Second Neumann eigenfunctions of the Laplace operator on $K_{n}$ have no interior critical points. In particular, the hot spots conjecture holds for these domains. Moreover, the second Neumann eigenvalue of these domains is simple.
We also use Proposition 1.5 to prove a more general result pertaining to the hot spots conjecture:
Theorem 1.7.
Let $\Omega$ be a convex Lipschitz domain, and let $u$ be a second Neumann eigenfunction of $\Omega$. Let $\Omega_{+}=\{x\in\Omega\mid u(x)>0\}$. If the diameter of $\Omega_{+}$ is at most half of the diameter of $\Omega$, then $u$ has no interior critical points in $\Omega_{+}$. The same statement applies to the set on which $u$ is negative.
2. First mixed eigenvalue limits
In this section, we prove an eigenvalue estimate that, together with Proposition 1.5, proves Theorem 1.1. We begin by recalling the variational characterization of the first mixed eigenvalue and eigenfunction. Let $H_{D}^{1}(\Omega)$ be the completion of the set of $C^{\infty}$ functions in $\overline{\Omega}$ that vanish on $D$ with respect to the norm
$\|u\|_{H_{D}^{1}}^{2}=\int_{\Omega}(|u|^{2}+|\nabla u|^{2}).$
Then we have
(2) $\lambda_{1}^{D}=\inf_{u\in H^{1}_{D}(\Omega)\setminus\{0\}}\frac{\int_{\Omega}|\nabla u|^{2}}{\int_{\Omega}|u|^{2}},$
and the minimizers of the functional above are exactly the first mixed eigenfunctions of the Laplacian. Constructing appropriate test functions for this functional allows us to give upper bounds on the first mixed eigenvalue.
By translating the domain $\Omega$, we may assume that the origin is contained in $D$. As in Section 1, let $d$ denote the diameter of $\Omega$:
$d=\text{diam}(\Omega)=\sup\{|x-y|:x,y\in\Omega\}.$
Let $V$ be the area of $\Omega\setminus B(0,d/4)$. Since $d$ is the diameter of $\Omega$, we have $V>0$.
Proposition 2.1.
Suppose that $D_{\epsilon}\subseteq B(0,\epsilon)$ for some $40<\epsilon<d/$ 4. Then we have
$\lambda_{1}^{D_{\epsilon}}\leq\frac{2\pi\ln(d/(4\epsilon))}{V\ln(4\epsilon/d)^{2}}.$
In particular, $\lambda_{1}^{D_{\epsilon}}$ tends to zero as $\epsilon$ tends to zero.
Proof.
For $40<\epsilon<d/$ 4, define a test function $\varphi$ in polar coordinates by
$\varphi_{\epsilon}(r)=\begin{cases}0\;\;&\text{if}\;\;r\leq\epsilon\\ 1-\frac{1}{\ln(4\epsilon/d)}\ln(4r/d)\;\;&\text{if}\;\;\epsilon<r<d/4\\ 1\;\;&\text{if}\;\;r\geq d/4.\end{cases}$
Note that $\varphi_{\epsilon}$ is an element of $H^{1}_{D_{\epsilon}}(\Omega)$ and is thus a valid test function for the variational formula for $\lambda_{1}^{D_{\epsilon}}$. We bound the Rayleigh quotient of $\varphi_{\epsilon}$ by bounding $\int_{\Omega}|\nabla\varphi_{\epsilon}|^{2}$ from above and $\int_{\Omega}\varphi_{\epsilon}^{2}$ from below. The quotient of these estimates will yield the desired estimate on $\lambda_{1}^{D_{\epsilon}}$. First,
[…]
$=\frac{2\pi\ln(d/(4\epsilon))}{\ln(4\epsilon/d)^{2}}.$
For the other bound, we simply use that $V>0$:
$\int_{\Omega}\varphi_{\epsilon}^{2}\geq\int_{\Omega\setminus B(0,d/4)}\varphi_{\epsilon}^{2}=V.$
Remark 2.2.
Our proof of Proposition 2.1 actually works without most of the restrictions on $\Omega$ and $D$ given in Section 1. For example, our proof applies to the case of any bounded Lipschitz domain $\Omega$. The proof also does not require $D$ to be connected or to be chosen such that eigenfunctions of (1) vanish pointwise on $D$.
Remark 2.3.
The fact that $\lambda_{1}^{D_{\epsilon}}$ tends to 00 as $\epsilon$ tends to zero is not new. One can find this fact, for instance, in [FNO21]. However, the authors in [FNO21] prove a significantly more precise result and state it in terms of the Sobolev capacity of $D$, so we chose to include a simple proof in this paper.
3. Proof of the main theorem
By Proposition 2.1, Theorem 1.1 follows from Proposition 1.5, which we prove in this section. We begin by reviewing some preparatory results on nodal sets (i.e. zero-level sets) of Laplace eigenfunctions.
Nodal sets of Laplace eigenfunctions have been studied for many years. It is well known that non-constant functions satisfying $-\Delta u=\lambda u$ on an open set have nodal sets that are unions of real-analytic arcs (see, e.g., [C76]) and, in particular, have no isolated points. If such $u$ vanishes at a critical point $p$, then there exists a disk neighborhood $U$ of $p$ such that $(U\cap u^{-1}(\{0\}))\setminus\{p\}$ is the union of at least four arcs. Moreover, $U$ can be chosen such that $u$ is positive in two connected components of $U\setminus u^{-1}(\{0\})$ and negative in the other two components. The following lemma places an additional restriction on the nodal sets of eigenfunctions whose eigenvalue is $\lambda_{1}^{D}$.
Lemma 3.1.
Let $u$ be a pointwise solution of $-\Delta u=\lambda_{1}^{D}u$ ($u$ is not necessarily equal to $u_{1}^{D}$) in $\Omega$ that extends continuously to $\overline{\Omega}$. Identify $u$ with its extension. If $u^{-1}(\{0\})$ contains a loop intersecting $\Omega$, then the region bounded by this loop intersects $D$.
Proof.
Recall the variational formulation (2) for $\lambda_{1}^{D}$ and the fact that the only minimizers of the functional are scalar multiples of $u_{1}^{D}$. Suppose for contradiction that $u^{-1}(\{0\})$ does contain a loop that bounds a topological disk $B$ that does not intersect $D$. Then $\varphi:=u\chi_{B}$ (where $\chi_{B}$ is the indicator function for $B$) is an element of $H^{1}_{D}(\Omega)$, and integration by parts gives
$\int_{\Omega}|\nabla\varphi|^{2}=\lambda_{1}^{D}\int_{\Omega}|\varphi|^{2}$
since $\varphi=0$ on $\partial\Omega$. This implies that $\varphi$ is a scalar multiple of $u_{1}^{D}$. Since the loop intersects $\Omega$, the set $\Omega\setminus B$ is non-empty and open, so we obtain a contradiction to unique continuation. ∎
Let $J_{0}$ be the order zero Bessel function of the first kind (see, for instance, Chapter 10 of [NIST]), and let $j_{0}\approx 2.4048$ be the first positive zero of $J_{0}(x)$. Note that $J_{0}(0)=1$, so $J_{0}$ is positive on $[0,j_{0})$. Note also that $J^{\prime}_{0}(0)=0$ and that $J^{\prime}_{0}(x)<0$ on $(0,j_{0}]$. Let $d$ be the diameter of $\Omega$, and suppose that $D$ is chosen such that $\lambda_{1}^{D}\leq\big{(}\frac{j_{0}}{d}\big{)}^{2}$.
Proof of Proposition 1.5.
Suppose toward a contradiction that $u_{1}^{D}$ has a critical point $p\in\Omega$. Since $u$ does not vanish anywhere in $\Omega\setminus D$, we may suppose without loss of generality that $u\geq 0$. By translating $\Omega$, we may suppose without loss of generality that $p$ equals the origin. Using polar coordinates, let $w(x,y)=J_{0}(\sqrt{\lambda_{1}^{D}}r)$. For $(x,y)\in\overline{\Omega}$, we have $r<d$ (with strict inequality since $p$ is an interior point), from which it follows that $w>0$ in $\overline{\Omega}$. By definition, $w$ satisfies $-\Delta w=\lambda_{1}^{D}w$ in $\Omega$. Define $f(x,y)=u_{1}^{D}(0)w(x,y)-u_{1}^{D}(x,y)$. This is a Laplace eigenfunction with eigenvalue $\lambda_{1}^{D}$. Moreover, $f$ is non-constant since $\lambda_{1}>0$ and $f>0$ on $D$. We will use $f$ to construct a test function that contradicts the variational formulation (2) for $\lambda_{1}^{D}$.
We have $f(0)=0$ and $\nabla f(0)=0$ since 00 is a critical point for both $u_{1}^{D}$ and $w$. Thus, there exists a neighborhood $U$ of 00 such that $U\setminus f^{-1}(\{0\})$ has at least four connected components, with $f$ taking alternating signs on each connected component. By Lemma 3.1 and using that $f>0$ on $D$, $\Omega\setminus f^{-1}(\{0\})$ has at least four connected components, and $f$ is positive in at least two of these components. Since $D$ is connected, there exists a unique connected component of $\Omega\setminus f^{-1}(\{0\})$ whose closure intersects $D$. Thus, there exists a connected component $B$ of $\Omega\setminus f^{-1}(\{0\})$ on which $f$ is positive and whose closure does not intersect $D$. Let $\varphi=f\chi_{B}\in H^{1}_{D}(\Omega)$.
We claim that $\partial_{\nu}\varphi<0$ almost everywhere on $\partial\Omega\cap\partial B$. Indeed, since $|(x,y)|\leq d$ on $\partial\Omega$ and since $u_{1}^{D}$ is Neumann on $\partial\Omega\setminus D$, the following holds for almost every $(x,y)\in\partial\Omega\setminus D$:
$\partial_{\nu}\varphi(x,y)=u_{1}^{D}(0)\partial_{\nu}w(x,y)=u_{1}^{D}(0)\sqrt{\lambda_{1}^{D}}J_{0}^{\prime}\Big{(}\sqrt{\lambda_{1}^{D}}|(x,y)|\Big{)}\frac{(x,y)}{|(x,y)|}\cdot\nu.$
Because $(\Omega,D)$ is Neumann convex, we have $(x,y)\cdot\nu\geq 0$ for all $(x,y)\in\partial\Omega\setminus D$, so $\partial_{\nu}\varphi\leq 0$ on $\partial\Omega\cap\partial B$. Using integration by parts and our knowledge of where $\varphi$ vanishes, we have
$\int_{\Omega}|\nabla\varphi|^{2}=\lambda_{1}^{D}\int_{\Omega}\varphi^{2}+\int_{\partial\Omega\cap\partial B}\varphi\partial_{\nu}\varphi\leq\lambda_{1}^{D}\int_{\Omega}\varphi^{2}.$
Dividing both sides of this inequality by $\int_{\Omega}\varphi^{2}$ gives a contradiction to the variational characterization (2) of $\lambda_{1}^{D}$ since $\varphi$ vanishes on a non-empty open set. ∎
Proof of Theorem 1.1.
By Proposition 1.5, we need only show that for sufficiently small $\epsilon$ and appropriate $D_{\epsilon}\subseteq B(0,\epsilon)$,
$\lambda_{1}^{D_{\epsilon}}\leq\Big{(}\frac{j_{0}}{d}\Big{)}^{2}.$
This follows from Proposition 2.1. ∎
4. Examples
In this section, we construct several examples that satisfy the hypotheses of Proposition 1.5 and Theorem 1.1 as well as examples illustrating the necessity of the connectivity hypothesis in these results. We begin with the latter.
Example 4.1.
Let $\Omega$ be the square given by the product of intervals $(-1,1)\times(-1,1)$. We first show that if we allow $D$ to have two connected components, then there exist arbitrarily small sets $D$ in $\Omega$ such that $u_{1}^{D}$ has an interior critical point. The second example shows that if we allow $D$ to have four connected components, then there exists arbitrarily small $D$ in $\Omega$ such that $u_{1}^{D}$ has an interior local maximum.
Let $0<\epsilon<\frac{1}{2}$. Let $D_{\epsilon}$ equal the union of the disks $B((\pm\epsilon,0),\epsilon/4)$. See Figure 1. Then the first eigenfunction $u_{1}^{D_{\epsilon}}$ is even about both the $x$- and $y$-axis since it is non-negative. Hence $\partial_{x}u_{1}^{D_{\epsilon}}\equiv 0$ on the $y$-axis and $\partial_{y}u_{1}^{D_{\epsilon}}\equiv 0$ on the $x$-axis. In particular, the origin is a critical point of $u$ for all such $\epsilon$.
Figure 1. The two pairs $(\Omega,D)$ constructed in Example 4.1 each with $\epsilon=0.1$. The dotted lines show the lines of symmetry for each figure. These graphics were created with Desmos (2024).
Now let $D_{\epsilon}$ be the union of the four disks $B((\pm\epsilon,\pm\epsilon),\epsilon/4)$ and $B((\pm\epsilon,\mp\epsilon),\epsilon/4)$. See Figure 1. Since $u$ is non-negative, simple, and even about the lines $y=\pm x$ and the $x$- and $y$-axes, $u$ restricts to a non-negative—and hence, first—mixed eigenfunction of the domain $P=\{(x,y)\mid 0<x<1,0<y<x\}\cap\Omega$ with Dirichlet conditions on the semi-circular arc in $\partial P$. By the argument given for Proposition 2.1 of [JM22], the vertex of $P$ at the origin is a local maximum for the restriction of $u_{1}^{D_{\epsilon}}$ to $P$. This vertex is then also an interior local maximum of $u$.
Theorem 4.2.
Let $P_{n}$ denote a regular $n$-gon with $n\geq 4$, and let $D$ be a line segment contained in an edge of $P_{n}$ ($D$ may equal the entire edge). Then
$\lambda_{1}^{D}\leq\Big{(}\frac{j_{0}}{d_{n}}\Big{)}^{2},$
where $d_{n}$ is the diameter of $P_{n}$. In particular, corresponding first mixed eigenfunctions have no interior critical points by Proposition 1.5.
Proof.
The case of $n=4$ is a simple computation that we omit (the below computation fails to give the estimate in this case). Suppose that $n\geq 5$. For ease of notation, we will temporarily suppress the subscript $n$ from $d_{n}$. Let $r=r_{n}$ denote the inradius of $P_{n}$ (that is, the supremum of radii of disks contained in $P_{n}$). Suppose that $D$ is contained in the $x$-axis and that $P_{n}$ is contained in the upper half-plane $\{(x,y)\mid y\geq 0\}$. Let $h$ be the length of the orthogonal projection of $P_{n}$ onto the $y$-axis, and let $\ell\geq h$. Let $\varphi(x,y)=\sin\Big{(}\frac{\pi}{2\ell}y\Big{)}$. Note that $\varphi$ is an appropriate test function for the variational characterization (2) of $\lambda_{1}^{D}$.
Let $B_{r}$ be the open disk of radius $r$ centered at $(0,r)$. Because $\varphi$ is increasing in $y$ for $0\leq y\leq h$, we get the lower bound
$\int_{P_{n}}\varphi^{2}\geq\int_{B_{r}}\varphi^{2}.$
We may, without loss of generality, suppose that $P_{n}$ is contained in the rectangle $R=[0,d]\times[0,\ell]$, from which we get the upper bound
$\int_{P_{n}}|\nabla\varphi|^{2}$ $\leq\int_{R}|\nabla\varphi|^{2}=\frac{d\pi^{2}}{8\ell}.$
Combining these estimates gives
(3) $\lambda_{1}^{D}\leq\frac{d\pi^{2}}{8\ell\int_{B_{r}}\sin^{2}\Big{(}\frac{\pi}{2\ell}y\Big{)}}.$
We can scale $P_{n}$ such that $d_{n}=1$. Let $\ell=1$. Then (3) gives the following bound that holds for all $n\geq 5$ (letting $r_{n}$ be the inradius of $P_{n}$):
$\lambda_{1}^{D}\leq\frac{\pi^{2}}{8\int_{B_{r}}\sin^{2}\Big{(}\frac{\pi}{2}y\Big{)}}.$
Since $r_{n}$ is increasing in $n$, we have $B_{r_{n}}\subseteq B_{r_{m}}$ for $n\leq m$, so the right-hand side of this inequality is decreasing in $n$. Hence, it suffices to prove that the right-hand side is less than $j_{0}^{2}$ for $n=5$. Numerical integration performed in Desmos (2024) gives
$\frac{\pi^{2}}{8\int_{B_{r_{5}}}\sin^{2}\Big{(}\frac{\pi}{2}y\Big{)}}<5.4<j_{0}^{2}.$
Proof of Theorem 1.6.
Let $u$ denote a second Neumann eigenfunction of $K_{n}$. Let $e$ be the line of reflection of $K_{n}$ as in the definition given in Section 1, and suppose that $e$ is contained in the $x$-axis. Then $u$ decomposes into a sum of even and odd eigenfunctions about this line:
$u_{\text{even}}(x,y)=\frac{1}{2}(u(x,y)+u(x,-y))$
$u_{\text{odd}}(x,y)=\frac{1}{2}(u(x,y)-u(x,-y)).$
We claim that $u_{\text{even}}$ is identically equal to zero. If not, then it restricts to a second Neumann eigenfunction of $P_{n}$. In this case, every second Neumann eigenfunction of $P_{n}$ extends via reflection to a second Neumann eigenfunction of $K_{n}$. By Lemma 3.3 of [JM20], the nodal set of $u_{\text{even}}$ cannot have distinct endpoints in $e$. Thus, the nodal set of $u_{\text{even}}$ has end points in two disjoint edges of $P_{n}$. By precomposing $u_{\text{even}}$ by an appropriate rotational symmetry, there exists another second Neumann eigenfunction $v$ of $P_{n}$ whose nodal set does not intersect $e$. Since the extension of $v$ to $K_{n}$ by reflection has three nodal domains, we get a contradiction to Courant’s nodal domains theorem, and the claim holds.
Thus, the second Neumann eigenspace of $K_{n}$ is spanned by extensions of first mixed eigenfunctions of $P_{n}$ with Dirichlet conditions on $e$. By the simplicity of the first mixed eigenvalue, the second Neumann eigenvalue of $K_{n}$ is simple. By the Hopf lemma, these eigenfunctions have no critical points in $e$. By Theorem 4.2, the theorem holds for $n\geq 4$. For $n=3$, the theorem is a special case of Theorem 1.1 of [H24a] or of Corollary 1.6 of [LY24]. ∎
Before proving Theorem 1.7, we remind the reader of an upper bound on the second Neumann eigenvalue on convex domains that, to the best our knowledge, first appeared in the literature as Corollary 2.1 in [BB99]. Their proof involves Brownian motion, so we provide a similar but analytic proof. Note that Bañuelos and Burdzy work with the normalization $-\frac{1}{2}\Delta$, so their estimate is stated with a “2” instead of a “4.”
Lemma 4.3.
Let $\Omega$ be a convex domain with diameter $d$. The second Neumann eigenvalue $\mu_{2}$ of the Laplacian on $\Omega$ is at most $4(j_{0}/d)^{2}$.
Proof.
Let $\ell$ be an open line segment contained in $\Omega$ with length $d$, and let $k$ denote the perpendicular bisector of $\ell$. Then $k$ segments $\Omega$ into two disjoint domains $\Omega_{\pm}$. By domain monotonicity, we have
$\mu_{2}\leq\max\{\lambda_{1}^{k}(\Omega_{-}),\lambda_{1}^{k}(\Omega_{+})\},$
so it suffices to prove that the desired upper bound holds for both of $\lambda_{1}^{k}(\Omega_{\pm})$. Let $x\in\partial\Omega_{+}$ be an endpoint of $\ell$. By translation, we may assume that $x$ is located at the origin. Then $B(0,\frac{d}{2})\cap\Omega$ is contained in $\Omega_{+}$. Define a test function $\varphi$ in polar coordinates by
$\varphi(r,\theta)=\begin{cases}J_{0}\Big{(}\frac{2j_{0}}{d}r\Big{)}\;\;&\text{if}\;\;(r,\theta)\in B(0,\frac{d}{2})\cap\Omega\\ 0\;\;&\text{otherwise}.\end{cases}$
Then $\varphi$ is a Laplace eigenfunction with eigenvalue $\Big{(}\frac{2j_{0}}{d}\Big{)}^{2}$ in $B(0,\frac{d}{2})\cap\Omega$ and is an element of $H^{1}_{k}(\Omega_{+})$. By the convexity of $\Omega$ and the non-negativity of $\varphi$, we have $\varphi\partial_{\nu}\varphi\leq 0$ on $\partial B(0,\frac{d}{2})\cap\partial\Omega_{+}$. Integration by parts then gives
$\int_{\Omega_{+}}|\nabla\varphi|^{2}=\int_{B(0,\frac{d}{2})\cap\Omega_{+}}|\nabla\varphi|^{2}=\Big{(}\frac{2j_{0}}{d}\Big{)}^{2}\int_{\Omega_{+}}\varphi^{2}+\int_{\partial B(0\frac{d}{2})\cap\partial\Omega_{+}}\varphi\partial_{\nu}\varphi\leq\Big{(}\frac{2j_{0}}{d}\Big{)}^{2}\int_{\Omega_{+}}\varphi^{2}.$
This proves that $\lambda_{1}^{k}(\Omega_{+})\leq 4\big{(}\frac{j_{0}}{d}\big{)}^{2}$, and the same argument applies to $\lambda_{1}^{k}(\Omega_{-})$. ∎
Proof of Theorem 1.7.
Let $\Omega$, $u$, and $\Omega_{+}$ be as in the theorem statement. Let $\mu_{2}$ be the eigenvalue associated to $u$. Because the restriction of $u$ to $\Omega_{+}$ is a first mixed eigenfunction with Dirichlet conditions on $u^{-1}(\{0\})$, we have $\lambda_{1}^{u^{-1}(\{0\})}(\Omega_{+})=\mu_{2}$. Let $d$ be the diameter of $\Omega$, and let $d_{+}$ be the diameter of $\Omega_{+}$. Since $d_{+}\leq\frac{1}{2}d$, we have
$\lambda_{1}^{u^{-1}(\{0\})}=\mu_{2}\leq 4\Big{(}\frac{j_{0}}{d}\Big{)}^{2}\leq 4\Big{(}\frac{j_{0}}{2d_{+}}\Big{)}^{2}=\Big{(}\frac{j_{0}}{d_{+}}\Big{)}^{2}$
by Lemma 4.3. By the discussion in Section 1, $(\Omega_{+},u^{-1}(\{0\}))$ is Neumann convex, so we may apply Proposition 1.5 to see that $u|_{\Omega_{+}}$ has no critical points in $\Omega_{+}$. To get the second statement of the theorem, we simply apply the same argument to the eigenfunction $-u$. ∎